Please help with my questions thanks in advance.
1) factor 112s^2+392sf+343f^2
2) f(x)=3x^2-15 need the x intercept
3) 4x(x-1)-5x(x)=3
4) x^1/6 * y^1/8 *z^1/3
Thanks a bunch

Answer by stanbon(75887)

1) factor 112s^2+392sf+343f^2
7*16s^2 + 7*56sf + 7*49f^2
---
7(16s^2 + 56sf + 49f^2)
---
7(4s + 7f)(4s + 7f)
=======================================

2) f(x)=3x^2-15 need the x intercept
Let y = 0 ; solve for "x"
3x^2 - 15 = 0
3(x^2 - 5) = 0
x = sqrt(5) or x = -sqrt(5)
========================================
3) 4x(x-1)-5x(x)=3
4x^2 - 4 - 5x^2 = 3
-x^2 = 7
x^2 = -7
x = sqrt(7)i or x = -sqrt(7)i
========================================

4) x^1/6 * y^1/8 *z^1/3
---
x^(4/24) * y^(3/24) * z(8/24)
---
(x^4 * y^3 * z*8)^(1/24)
================================
Cheers,
Stan H.